OpenStudy (anonymous):

what is the solution to the following system? -3x-2y=-12?.......... i need help fast,Please and thank you.Also an explanation on how you solved it would be much appreciated :)

OpenStudy (anonymous):

Is there more to the system than just that?

OpenStudy (anonymous):

oh gosh ,im so sorry.... the whole thing 9y+6x=-9 , -3x-2y=-12. @PlatypusFeet

OpenStudy (anonymous):

get y alone in each equation and then set them equal to eachother

OpenStudy (anonymous):

how would you do that? @Costanza

OpenStudy (anonymous):

Alright, let's try solving this using substitution. First let's take the second equation -3x-2y=-12 and get y by itself. -2y=3x-12 y=(-3/2)x+6 Now since we know what y is, we can plug it into the other equation. 9((-3/2)x+6)+6x=-9 Simplifies to - (-27/2)x+54+6x=-9 (-15/2)x=-63 x=8.4 To find y just plug x back into an equation. -3(8.4)-2y=-12 Simplifies to- y = -6.6 So x = 8.4 and y = -6.6 Just plug these back in to check and see if they're correct. 9(-6.6)+6(8.4)=-9 -> -59.4+50.4=-9 -> -9=-9 -3(8.4)-2(-6.6)=-12 -> -25.2 + 13.2 = -12 -> -12=-12

OpenStudy (anonymous):

\[9y+6x=-9 \\ -3x-2y=-12\] multiply the second one by \(3\) and you get \[9y+6x=-9 \\-9x-6y=-36\]

OpenStudy (anonymous):

add and you get \(0=-45\) which means the lines are parallel and do not intersect

OpenStudy (anonymous):

you can kind of see they are parallel from the start since the ratio of the coefficients are the same

OpenStudy (anonymous):

satellite you can't cancel out 6x with -6y

OpenStudy (anonymous):

maybe i read it wrong

OpenStudy (anonymous):

you can do it your way if you multiply the second one by 2 to get 9y+6x=-9 -6x-4y=-24

OpenStudy (anonymous):

oooh i made a mistake, sorry

OpenStudy (anonymous):

then 5y=-33 y=-6.6

OpenStudy (anonymous):

yeah lol its okay

OpenStudy (anonymous):

Thank you so much to all of you!!

OpenStudy (anonymous):

\[6x+9y=-9 \\-6x-4y=-24\]

OpenStudy (anonymous):

The elimination method would be faster had I been thinking.

OpenStudy (anonymous):

but wait... so my answer choices were 2,1, -2,-1 i infinate solutions and no solutions...

OpenStudy (anonymous):

does this mean it would be infinate?

OpenStudy (anonymous):

those are the only options?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

@PlatypusFeet @satellite73

OpenStudy (anonymous):

The answer shouldn't be any of those. Is it possible the equations were written down wrong earlier?

OpenStudy (anonymous):

idk what to say. It's not 2,1 -2,-1 or no solutions cause we found a solution but i dont think there are an infinite amount of solutions

OpenStudy (anonymous):

Infinite solutions would mean that you can plug any number into the equations and it would work.

OpenStudy (anonymous):

9y+6x=-9 , -3x-2y=-12 Those were the two equations.. it says solve by elimination? i probably should've mentioned that .so sorry

OpenStudy (anonymous):

@PlatypusFeet , @Costanza , @satellite73

OpenStudy (anonymous):

elimination is the method that satellite and i used earlier

OpenStudy (anonymous):

To solve through elimination you would multiply the second equation by 2: -6x-4y=-24 Add with first equation: 6x+9y=-9 -6x-4y=-24 You should get 5y=-33, which means y is equal to -6.6 Solving through elimination doesn't change the answer.

OpenStudy (anonymous):

Oh okay good..Thats whats confusing me though,these damned answer choices..let me double check to make sure i got this equation right...

OpenStudy (anonymous):

alright im sorry im an idiot. i put 9y+6x=-9 , -3x-2y=-12 .... it should be -3x-2y=-12.....9x+6y=-9 @PlatypusFeet @Costanza

OpenStudy (anonymous):

that's alot easier lol

OpenStudy (anonymous):

Then its no solution.

OpenStudy (anonymous):

i accidently switched the x & ys while retyping..sorry about that lol..

OpenStudy (anonymous):

no solution? alrite thank you guys so much!!! such a big help :)

OpenStudy (anonymous):

@satellite73 explained it earlier why they cancel each other out, making 0 = -45, which is a false answer. This makes it no solution.

OpenStudy (anonymous):

oh okay i see.... um.. care to help with one more equation?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

x-y=11 ****-x+y=-11 ?

OpenStudy (anonymous):

3,4 -3,-4 no solutions infinite solutions

OpenStudy (anonymous):

I believe that would be infinite solutions. x-y=11 -x+y=-11 Added together gets: 0=0 Meaning any answer could work.

OpenStudy (anonymous):

Your a life saver... thankyou again..